3/14x - 1/14y = 1/2 and 6x = 2(y+5)
The directions state to describe each system of equations as consistent and independent, consistent and dependent, or inconsistent

Answer by rapaljer(4671)

3/14x - 1/14y = 1/2
Multiply both sides by the LCD which is 14:
14(3/14x - 1/14y) = 14*1/2
3x - y = 7

For the second equation, you need to put it in standard form. Remove parentheses first:
6x = 2y+10
6x - 2y=10

Divide both sides by 2:
3x-y=5

The system to be solved is
3x-y=7
3x-y=5

Now, subtract the second equation from the first equation:
3x-y-(3x-y)=7-5
3x-y-3x+y=2
0=2

NO WAY!! Since there is NO SOLUTION, this means that the lines are PARALLEL, and the system is said to be "INCONSISTENT."

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus